OpenStudy (pulsified333):

use the following equation for the demand curve: q=400−4p (a) Enter the equation for revenue R as a function of price p : R= (b) What is the (instantaneous) rate of change of revenue with respect to price when the price is $ 13? (c) At what price is the derivative of revenue with respect to price exactly zero?

OpenStudy (pulsified333):

@mathmale

OpenStudy (mathmale):

Hope I can actually be of help; haven't worked with this type of algebra problem for some years now. But anyway: Revenue equals price times quantity. If I sell you 50 pens for $2 each, the revenue coming my way would be (50 pens * $2.00/pen), or $100. Thus if you have a formula relating price and quantity, which you do in q=400−4p, just solve this for p (price) and multiply the result by q (quantity). Read this carefully , please and let me know how much sense this makes to you before you get involved in calculations.

OpenStudy (mathmale):

You want to express revenue in terms of price, p, which means eliminating q. Are you able to do that algebraically?

OpenStudy (pulsified333):

i'm not sure

OpenStudy (mathmale):

I am contradicting myself, I'm afraid. We want to eliminate quantity (q), and keep price (p), so we must use q=400−4p; in other words, substitute 400-4p for quantity, q. Again, Revenue is price times quantity; here it's R(p)=p(400-4p). Expand that, please.

OpenStudy (kayders1997):

I just got done doing this! I could probably help

OpenStudy (pulsified333):

okay R(p)=400p-4p^2

OpenStudy (mathmale):

@kayders1997: Thank you for your offer, but at the moment Pulsified333 and I are doing fine.

OpenStudy (kayders1997):

I just say when you said I haven't done this in a long time sorry

OpenStudy (kayders1997):

saw

OpenStudy (mathmale):

Next part: you want the instantaneous rate of change of R with respect to price. When you see "instantaneous rate of change," that should remind you of a sophisticated technique we should apply next. What is that technique?

OpenStudy (pulsified333):

is it taking the derivative?

OpenStudy (mathmale):

Exactly. Please find dR/dp, the inst. rate of change of Revenue with respect to price.

OpenStudy (pulsified333):

R'(p)=400-8p

OpenStudy (pulsified333):

then i plug in 13 for P right?

OpenStudy (mathmale):

That is the general formula for the inst. r. of c. The probelm statement gives you a certain value, $13, for price, p. Yes. plug in $13 and evalute the "inst. r. of c. of Rev. when p=$13."

OpenStudy (pulsified333):

R'(p)=400-8p, then R'(13)=400-8(13)=296

OpenStudy (mathmale):

Great. But let's not forget the units of measurement. What are you measuring here?

OpenStudy (pulsified333):

$296?

OpenStudy (mathmale):

Partially right. But "inst. rate of change" implies that TWO different quantities are changing, not just $. What's the other quantity? Hint: What is our independent variable in "inst. rate of change" here?

OpenStudy (pulsified333):

400

OpenStudy (pulsified333):

I don't think the question actual gives us units to work with

OpenStudy (mathmale):

See if you can find your cue here: "What is the (instantaneous) rate of change of revenue with respect to price when the price"

OpenStudy (pulsified333):

@mathmale Revenue is the other quantity?

OpenStudy (mathmale):

You're looking to find the inst. rate of change of Revenue (dependent variable) with respect to price.

OpenStudy (pulsified333):

oh

OpenStudy (pulsified333):

i am still confused on what the dependent variable is ;(

OpenStudy (mathmale):

Revenue depends upon price.

OpenStudy (pulsified333):

so i was right?

OpenStudy (mathmale):

Express your result, for the inst. rate of ch. of revenue with respect to price, when price is $13, as $296 per dollar increase in price.

OpenStudy (pulsified333):

oh okay

OpenStudy (pulsified333):

onto the last part? which i think i know how to do

OpenStudy (pulsified333):

0=400-8p, then solve for P which P=50

OpenStudy (mathmale):

Great working with you. Let me know if you have further questions. I'm hitting the sack now, however. Very good. Be sure not to forget units of measurement.

OpenStudy (pulsified333):

I wont :D thank you

OpenStudy (mathmale):

My pleasure! Good night.

OpenStudy (pulsified333):

Night